# Kandinsky 3.0 Technical Report

### v2

@@ Abstract
$*{}^{*}$orresponding authors: Andrey Kuznetsov $<$kuznetsov@airi.net$>$, Denis Dimitrov $<$dimitrov@airi.net$>$. present Kandinsky ${}^{**}$he family of models is named after Wassily Kandinsky, the great Russian artist and an art theorist, the father of abstract art., a large-scale text-to-image generation model based on latent diffusion, continuing the series of text-to-image Kandinsky models and reflecting our progress to achieve higher quality and realism of image generation. Compared to previous versions of Kandinsky 2.x, Kandinsky 3.0 leverages a two times larger U-Net[1] backbone, a ten times larger text encoder and removes diffusion mapping. We describe the architecture of the model, the data collection procedure, the training technique, and the production system of user interaction. We focus on the key components that, as we have identified as a result of a large number of experiments, had the most significant impact on improving the quality of our model compared to the others. By our side-by-side comparisons, Kandinsky becomes better in text understanding and works better on specific domains. The project is available at https://ai-forever.github.io/Kandinsky-3.
@@ 1 Introduction
This paper presents Kandinsky 3.0, a new text-to-image generation model based on latent diffusion [6]. Earlier, we introduced other models of the Kandinsky family [12], the architecture of which is based on a two-stage pipeline using Diffusion Mapping between elements of latent vector spaces of images and text with subsequent decoding. In the Kandinsky 3.0 model, we focused on improving the text understanding, the image quality and simplifying the architecture by providing a single-stage pipeline in which generation takes place directly using text embeddings without any additional priors. The whole pipeline contains 11.9 billion parameters, almost three times more than the largest of the previous models of the Kandinsky family. Also, we integrated Kandinsky 3.0 into our user-friendly interaction system. We made our model completely public to promote the development of new technologies and openness in the scientific community.
@@
  • Firstly, we describe the user interaction demo system;
  • Secondly, we describe the key components of the Kandinsky 3.0 model, datasets usage strategy, training techniques, and various applications;
  • Finally, we report the results of side-by-side comparisons based on human preferences and discuss the model limitations.
2 Demo System
@@
We also created a Telegram-bot [^3] https://t.me/kandinsky21_bot in which text-to-image generation is available. The user can choose a model version (Kandinsky 2.1, 2.2, or Kandinsky 3.0) and view the current negative prompt or specify a new one.
@@ 3.1 Overall pipeline
Kandinsky 3.0 is a latent diffusion model, the whole pipeline of which includes a text encoder for processing a prompt from the user, a U-Net for predicting noise during the denoising (reverse) process, and a decoder for image reconstruction from the generated latent (Fig. 1). The text encoder and image decoder were completely frozen during the U-Net training. The whole model contains 11.9 billion parameters. Table 1 shows the number of parameters for the components of the Kandinsky 3.0 model and the Kandinsky 2.2 [12] and SDXL [10] models. Below, we take a look at each component of our new model.
@@ 3.2 U-Net architecture
  • At the initial stage, in high resolution, we decided to process the image using only convolutional blocks, while more compressed representations were fed to the transformer layers too. This ensures the global interaction of image elements.
We also reviewed the MaxViT architecture, which is almost entirely based on transformer blocks adapted to work with images by reducing the quadratic complexity of self-attention [20]. In the classification task, this architecture shows the best results in comparison with the models mentioned above. Despite this, during experiments, we found out that this architecture does not show good results in the generation task.
@@ 3.3 Text encoder
For the text encoder, we use the 8.6B encoder of the Flan-UL2 20B model [25], which is the most powerful encoder-decoder open-source model for now and is based on the pre-trained UL2 20B [26]. In addition to pretraining on a large corpus of texts, Flan-UL2 was also trained using supervised finetuning on many language tasks using Flan Prompting [27]. Our experiments showed that such finetuning also significantly improves image generation.
@@ 3.4 Sber-MoVQGAN
where $F^{i-1}$ is the intermediate feature map, $\mu$ and $\sigma$ are the functions for calculating the mean and standard deviation of the activation, $\phi_{\gamma}$ and $\phi_{\beta}$ are the trainable affine transformations, which convert $z_{q}$ into the scaling and bias values. Other important features of our implementation include the addition of EMA (exponential moving average) weights and a modified loss function from ViT-VQGAN [31] during the training stage.
@@
Table 2: Sber-MoVQGAN comparison with competitors on ImageNet dataset.
Model Latent size Num Z Train steps FID normal-$\downarrow$ SSIM normal-$\uparrow$ PSNR normal-$\uparrow$ L1 normal-$\downarrow$
@@ 4 Data and Training Strategy
We also divided all the data into two categories. We used the first at the initial stages of low-resolution pretraining and the second for mixed and high-resolution finetuning at the last stage. The first category includes open large text-image datasets such as LAION-5B [36] and COYO-700M [37] and “dirty” data that we collected from the Internet. The second category contains the same datasets but with stricter filters, especially for the image aesthetics quality.
@@ 5.1 Inpainting and Outpainting
Implementation of the inpainting model is the same as GLIDE [38]: we initialize our model from base Kandinsky model weights. Then, we modify the input convolution layer of U-Net (Sec. 3.2) so that the input can additionally accept the image latent and mask. Thus, U-Net takes as many as 9 channels as input: 4 for the original latent, 4 for the image latent, and an additional channel for the mask. We zeroed the additional weights, so training starts with the base model.
For training, we generate random masks of the following forms: rectangular, circles, strokes, and arbitrary form. For every image sample, we use up to 3 masks, and for every image, we use unique masks. We use the same dataset as for the training base model with generated masks. We train our model using Lion [39] with lr=1e-5 and apply linear warmup for the first 10 thousand steps. We train our model for 250 thousand steps. The inpainting results are in Fig. 3. The outpainting results can be found in the Outpainting appendix section.
@@ 5.2 Image-to-Video Generation
The scene generation process involves defining depth values along the $z$-axis within the interval $[(z_{\text{near}},z_{\text{far}})]$ in the coordinate system. Depth estimation utilizes either MiDaS [41] or AdaBins [42]. The camera’s first-person motion trajectory, simulated through functions $x(t)$, $y(t)$, $z(t)$, $\alpha(t)$, $\beta(t)$, and $\gamma(t)$, includes perspective projection operations with the camera initially fixed at the origin and the scene at a distance of $z_{\text{near}}$. Then, we apply transformations by rotating points around axes passing through the scene’s center and translating to this center. Acknowledging the limitations of a single-image-derived depth map, addressing distortions resulting from minor camera orientation deviations is crucial. Two essential considerations follow: adjusting scene position through infinitesimal transformations and employing the image-to-image approach after each transformation. The image-to-image technique facilitates the realization of seamless and semantically accurate transitions between frames, enhancing the aesthetic appeal of this approach. The examples of image-to-video generations are presented in the Image-to-Video appendix section.
@@ 5.3 Text-to-Video
Based on the Kandinsky 3.0 model, we also created the text-to-video generation pipeline Kandinsky Video [43], consisting, in addition to a text encoder and an image decoder, of two models – for keyframes generation and interpolation between them. Both models use the pretrained weights of Kandinsky 3.0 as the backbone. We have also implemented the publicly available text-to-video generation interface [^6] https://fusionbrain.ai/en/t2v/ in the Fusionbrain.ai website, which we mentioned above (Sec. 2). Please refer to the main paper for additional details regarding the text-to-video model.
@@ 6 Human evaluation
Some examples of text-to-image generations can be found in the Text-to-Image appendix section. To compare Kandinsky 3.0 with other well-known models, we have compiled a balanced set of 2.1K prompts in 21 categories. Using this set, we first performed several side-by-side (SBS) comparisons between different versions of the Kandinsky 3.0 and then selected the best version. We conducted three rounds of SBS comparisons involving 28 people to do this. Next, we conducted side-by-side comparisons of Kandinsky 3.0 with the Kandinsky 2.2 [12], SDXL [10] and DALL-E 3 [11] models. Each study involved 12 people who voted an average of 30,000 times in each SBS. For this purpose, we developed chatbot that showed one of 2.1K pairs of images.
@@
We compared the visual quality and text comprehension in total for all categories, and each category separately. The visual examples of Comparison to prior works and Human evaluation results are presented in the appendix.

### v3

@@ Abstract
WeCorresponding authors: Andrey Kuznetsov $<$kuznetsov@airi.net$>$, Denis Dimitrov $<$dimitrov@airi.net$>$. present Kandinsky 3.0∗∗The family of models is named after Wassily Kandinsky, the great Russian artist and an art theorist, the father of abstract art., a large-scale text-to-image generation model based on latent diffusion, continuing the series of text-to-image Kandinsky models and reflecting our progress to achieve higher quality and realism of image generation. In this report we describe the architecture of the model, the data collection procedure, the training technique, and the production system for user interaction. We focus on the key components that, as we have identified as a result of a large number of experiments, had the most significant impact on improving the quality of our model compared to the others. We also describe extensions and applications of our model, including super resolution, inpainting, image editing, image-to-video generation, and a distilled version of Kandinsky 3.0 – Kandinsky 3.1, which does inference in 4 steps of the reverse process and 20 times faster without visual quality decrease. By side-by-side human preferences comparison, Kandinsky becomes better in text understanding and works better on specific domains. The code is available at https://github.com/ai-forever/Kandinsky-3.
@@ 1 Introduction
This paper presents Kandinsky 3.0, a new text-to-image generation model based on latent diffusion [5]. Earlier, we introduced other models of the Kandinsky family [11], the architecture of which is based on a two-stage pipeline using Diffusion Mapping between elements of latent vector spaces of images and text with subsequent decoding. In the Kandinsky 3.0 model, we focused on improving the text understanding, the image quality and simplifying the architecture by providing a single-stage pipeline in which generation takes place directly using text embeddings without any additional priors. The whole pipeline contains 11.9 billion parameters, almost three times more than the largest of the previous models of the Kandinsky family. Also, we integrated Kandinsky 3.0 into our user-friendly website interaction system. We made our model completely public to promote the development of new technologies and openness in the scientific community.
@@
  • Firstly, we describe the user interaction system (Section 2);
  • Secondly, we describe the key components of the Kandinsky 3.0 model architecture (Section 3), dataset usage strategy and training techniques (Section 4);
  • We also report additional features such as distillation (aka Kandinsky 3.1, see Section 5.1), super resolution (aka Kandinsky SuperRes, Section 5.2) and prompt beautification (Section 5.3), and various applications: inpainting (Section 6.1), image editing (Section 6.2), image-to-video (Section 6.3) and text-to-video (Section 6.4);
  • Finally, we describe human evaluation methodology (Section 7) and report the extensive results of side-by-side comparisons based on human preferences (see in appendix C. Side-by-side human evaluation results).
2 Interaction System
@@
We also created a Telegram-bot [^3] https://t.me/kandinsky21_bot in which text-to-image generation is available.
@@ 3.1 Overall pipeline
Kandinsky 3.0 is a latent diffusion model, the whole pipeline of which includes a text encoder for processing a prompt from the user, a U-Net [12] for predicting noise during the denoising (reverse) process, and a decoder for image reconstruction from the generated latent (Fig. 1). The text encoder and image decoder were frozen during the U-Net training. The whole model contains 11.9 billion parameters. Table 1 shows the number of parameters for the components of the Kandinsky 3.0 model, the Kandinsky 2.2 [11] and SDXL [9] models. Below, we take a look at each component of our new model.
@@ 3.2 U-Net architecture
  • At the first network layers, in high resolution level, we decided to process the latents using only convolutional blocks, while more compressed latent representations were fed to the transformer layers too. This ensures the global interaction of image elements.
We also reviewed the MaxViT architecture [20], which is almost entirely based on transformer blocks adapted to work with images by reducing the quadratic complexity of self-attention. In the classification task, this architecture shows the best results in comparison with the models mentioned above. Despite this, during experiments, we found out that this architecture does not show good results in the generation task.
@@ 3.3 Text encoder
For the text encoder, we use the 8.6B encoder of the Flan-UL2 20B model [25], which is based on the pre-trained UL2 20B [26]. In addition to pretraining on a large corpus of texts, Flan-UL2 was also trained using supervised fine-tuning on many language tasks using Flan Prompting [27]. Our experiments showed that such fine-tuning also significantly improves image generation.
@@ 3.4 Sber-MoVQGAN
where $z_{q}$ is the latent consisting of codebook codes, $F^{i-1}$ is the intermediate feature map, $\mu$ and $\sigma$ are the functions for calculating the mean and standard deviation of the activation, $\phi_{\gamma}$ and $\phi_{\beta}$ are the trainable affine transformations, which convert $z_{q}$ into the scaling and bias values. Other important features of our implementation include the addition of EMA (exponential moving average) weights and a modified loss function from ViT-VQGAN [31] during the training stage.
@@
Table 2: Sber-MoVQGAN comparison with competitors on ImageNet dataset. Our 270B version outperform all other models in terms of all quality metrics.
Model Latent size Num Z Train steps FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ L1 $\downarrow$
@@ 4 Data and Training Strategy
We also divided all the data into two categories. We used the first at the initial stages of low-resolution pretraining and the second for mixed and high-resolution fine-tuning at the last stage. The first category includes open large text-image datasets such as LAION-5B [36] and COYO-700M [37] and “dirty” data that we collected from the Internet. The second category contains the same datasets but with stricter filters, especially for the image aesthetics quality.
@@
5 Additional Features
5.1 Kandinsky 3.1 (aka Kandinsky 3.0 Distilled)
Figure 3: Discriminator architecture for Kandinsky 3.1. Gray blocks inherit the weight of Kandinsky 3.0 and remain frozen during training.
A serious problem of diffusion models is the generation speed. To obtain a single image, it is usually necessary to go through several dozen steps (for example, 50) in the reverse process, that is, to pass data through U-Net many times with a batch size of 2 (with conditioning and without it) for classifier free guidance. To solve this problem, we used the Adversarial Diffusion Distillation approach [38], but with a number of significant modifications:
  1. 1.
    If pretrained pixel models were used as a discriminator, it would be necessary to decode the generated image using MoVQ and throw gradients through it, which would lead to huge memory costs. These costs do not allow training the model in $1024\times 1024$ resolution. Therefore, as a discriminator, we used the frozen downsample part of the U-Net from Kandinsky 3.0 with trainable heads after each layer of resolution reduction (Fig. 3);
  2. 2.
    We added cross-attention on text embeddings from FLAN-UL2 to the discriminator heads instead of adding text CLIP-embeddings. This improved the text alignment using a distilled model;
  3. 3.
    We used Wasserstein Loss [39]. Unlike Hinge Loss, it is unsaturated, which avoids the problem of zeroing gradients at the first stages of training, when the discriminator is stronger than the generator;
  4. 4.
    We removed the regularization in the Distillation Loss, since according to our experiments it did not affect the quality of the model;
  5. 5.
    We found that the generator quickly becomes more powerful than the discriminator, which leads to learning instability. To solve this problem, we have significantly increased the learning rate of the discriminator. For the discriminator, we set the learning rate is equal to $31e-$ 3, and for the generator $51e-$ 5. To prevent divergence, we also used gradient penalty, as in the [38].
We trained a distilled model on a dataset with 100 thousand of highly-aesthetic image-text pairs, which we manually selected from the pretraining dataset. As a result, we speed up the Kandinsky 3.0 by almost 20 times, making it possible to generate an image in only 4 passes through U-Net. The acceleration is also due to the fact that there is no need to use classifier free guidance in the distilled version. However, like in [38], for a serious acceleration, we had to sacrifice the quality of the text comprehension, which can be seen from the human evaluation results of side-by-side comparison (Section Human evaluation results for distillation). Generation examples by Kandinsky 3.1 can be found in the sections Distillation and Comparison to prior works.
5.2 Kandinsky SuperRes
Figure 4: Comparison of Kandinsky SuperRes, Stable Diffusion and Real-ESRGAN models at 1024 resolution. Better to zoom image.
Based on Kandinsky 3.0, we developed the Kandinsky SuperRes model to generate high-resolution 4k images. Here we describe the modifications that we made.
  1. 1.
    For Kandinsky SuperRes model we used diffusion in pixel space instead of latent diffusion. This eliminated the loss of quality when encoding and decoding images using an autoencoder. In addition, our experiments have shown that model based on pixel diffusion converges faster and better in this task than with latent diffusion.
  2. 2.
    We implemented Efficient U-Net, similar to the one described in the Imagen [6]. Compared to U-Net from Kandinsky 3.0, Efficient U-Net consumes less memory and also has better convergence. Instead of 3 residual blocks at each downscaling, Efficient U-Net uses more low-resolution blocks and fewer high-resolution blocks. The order of convolution and downsampling/upsampling operations is changed relative to the original U-Net. In addition, we removed conditioning for text prompt, because it does not contribute to 4K high resolution generation. As a result, Efficient U-Net of Kandinsky SuperRes contains 413M parameters.
  3. 3.
    During training, Efficient U-Net predict $x_{0}$ (i.e., the original image) instead of the noise level at a given time t, which avoided problems with changing the color of the generated SR image.
The training was carried out in 2 stages. First, the model learned on the LAION dataset [36] on 32 A100 for 1570 thousand steps with a batch size of 2 for a resolution of 256 to 1024. Then we trained the model on the aesthetic high-resolution sets, which we used to train Kandinsky 3.0 for 1500 thousand steps. At the second stage of training, we included JPEG compression similar to the scheme in the Real-ESRGAN [40].
The Kandinsky SuperRes model can work with images of various resolutions, but the main goal is 4K high-resolution generation. Since the Kandinsky SuperRes model was already trained for resolutions from 256 to 1024, higher resolution training was not possible due to GPU Tesla A100 memory overflow. For this reason, we used the MultiDiffusion algorithm [41] to generate 4k images. More specifically, first we divide the image into overlapping patches, and then at each diffusion step we remove the noise and average the pixel/latent values of the overlapping areas. Thus, having gone through all the stages of diffusion, we obtain a seamless image of any resolution. Kandinsky SuperRes model works in inference in 5 steps using DPMSolverMultistepScheduler [^6] https://huggingface.co/docs/diffusers/api/schedulers/multistep_dpm_solver. The inference time for image generation in 4K resolution takes 13 seconds, and in 1K it takes 0.5 seconds.
The table 3 shows a comparison of Kandinsky SuperRes with the Real-ESRGAN  [^7] https://github.com/ai-forever/Real-ESRGAN/ and Stable Diffusion x4 Upscaler  [^8] https://huggingface.co/stabilityai/stable-diffusion-x4-upscaler [5] models in terms of FID, SSIM, PSNR and L1 metrics on the our own dataset Wikidata 5K and RealSR(V3)  [^9] https://github.com/csjcai/RealSR/tree/master [42] and Set14  [^10] https://paperswithcode.com/dataset/set14 [43] datasets. Our own dataset Wikidata 5K contains 5 thousand images collected from Wikipedia in 1K resolution. RealSR(V3) contains 100 test images in 1K and 2K resolutions. Set14 contains 14 low resolution images with JPEG artifacts. As a result, the Kandinsky SuperRes model shows the best results in terms of all quality metrics for all datasets. Figure 4 shows examples of generation of Kandinsky SuperRes, Stable Diffusion and Real-ESRGAN models at a resolution of 1024. Figure 5 shows examples of generation of Kandinsky SuperRes in 4K resolution. We released the weights and code for Kandinsky SuperRes under an open-source license  [^11] https://github.com/ai-forever/KandiSuperRes/.
Figure 5: Example of Kandinsky SuperRes generation in 4K resolution. Better to zoom image.
Table 3: Comparison of Kandinsky SuperRes, Real-ESRGAN [40] and Stable Diffusion [5] models. Kandinsky SuperRes outperforms other models in terms of all quality metrics for all test datasets.
Datasets Model FID↓ PSNR↑ SSIM↑ L1↓
Real-ESRGAN 9.96 24.48 0.73 0.0428
Stable Diffusion 3.04 25.05 0.67 0.0435
Wikidata 5k Kandinsky SuperRes 0.89 28.52 0.81 0.0257
Real-ESRGAN 73.26 23.12 0.72 0.0610
Stable Diffusion 47.79 24.85 0.67 0.0493
RealSR(V3)[42] Kandinsky SuperRes 47.37 25.05 0.75 0.0462
Real-ESRGAN 115.94 22.88 0.62 0.0561
Stable Diffusion 76.32 23.60 0.57 0.0520
Set14 [43] Kandinsky SuperRes 61.00 25.70 0.70 0.0390
5.3 Prompt beautification
Many diffusion text-to-image models have some kind of inconvenience due to the fact that the visual quality and detail of the generation depends on the degree of detail of the text prompt. Sometimes, in practice, user has to use long, redundant prompts to generate desirable images. Normal user prompts are not such. To solve this problem, we have built a function into the generation pipeline to add details to the user’s prompt using LLM. An instruction is sent to the input of the language model with a request to improve the prompt, and the model’s response is sent as the input into Kandinsky 3.0 model. As an LLM, we used Neural-Chat-7b-v3-1 [44] (based on Mistral 7B [45]) with the following system instruction:
[…]
Here {prompt} is the user’s text prompt. Examples of generation for the same prompt with and without beautification are presented in the section Prompt beautification. We conducted a side-by-side human evaluation comparison of the generation quality using the prompt beautification and without it. We conducted comparison for both Kandinsky 3.0 and Kandinsky 3.1 to assess how strongly the language model affects the generated images. Each generation was evaluated by visual quality and by the text-image alignment. The results of the comparison can be seen in the section Human evaluation results for prompt beautification. In general, human preferences are definitely more inclined towards generations with prompt beautification. The only exception is the correspondence of the generated image to the text for the Kandinsky 3.1 model. This is due to the fact that the distilled model generally understands the text worse.
@@ 6.1 Inpainting and Outpainting
Implementation of the inpainting model is the same as GLIDE [46]: we initialize our model from base Kandinsky model weights. Then, we modify the input convolution layer of U-Net (Section 3.2) so that the input can additionally accept the image latent and mask. Thus, U-Net takes as many as 9 channels as input: 4 for the original latent, 4 for the image latent, and an additional channel for the mask. We zeroed the additional weights, so training starts with the base model.
For training, we generate random masks of the following forms: rectangular, circles, strokes, and arbitrary form. For every image sample, we use up to 3 masks, and for every image, we use unique masks. We use the same dataset as for the training base model with generated masks. We train our model using Lion [47] with lr=1e-5 and apply linear warmup for the first 10k steps. We train our model for 250 thousand steps. The inpainting results are in Fig. 6. The outpainting results can be found in the Outpainting appendix section. Additionally, we finetune our model using object detection datasets and LLaVA captions for 50k steps.
@@
6.2 Image editing
Kandinsky 3.0 provides image generation not only using a text prompt, but also using an image as a visual prompt. This allows one to edit an existing image, change its style and add new objects to it. To do this, we extended an IP Adapter-based approach [48]. To implement our own IP-Adapter based on our basic generation model, we used attention adapters. We used the ViT-L-14 as the CLIP model [15]. We get CLIP-embeddings of the size batch size $\times$ 768, which are then transformed by a linear layer into tensors of the size batch size $\times$ 4 $\times$ 4096. By adding a couple of new layers for key and value images in the cross-attention mechanism, we sum up the output of a text cross-attention with the output of cross-attention for images. We trained the IP-Adapter on the COYO 700m dataset [37] with a batch size 288 during 800 thousand iterations. As a result, Kandinsky 3.0 supports image variation, image-image mixing, and image-text mixing. For generation examples see section Image Editing with IP-Adapter.
In addition, we found that the IP Adapter-based approach does not preserve the shape of objects in the image, so we decided to train ControlNet [49] in addition to our generation model to consistently change the appearance of the image, preserving more information compared to the original one. We used HED detector [50] as a model to obtain the edges in the image fed to the ControlNet input. The training lasted 5 thousand iterations on the COYO 700m dataset [37] on 8 Tesla A 100 GPU with a batch size 512.
@@ 6.3 Image-to-Video Generation
The scene generation process involves defining depth values along the $z$-axis within the interval $[(z_{\text{near}},z_{\text{far}})]$ in the coordinate system. Depth estimation utilizes either MiDaS [52] or AdaBins [53]. The camera is characterized by the coordinates $(x,y,z)$ in the three-dimensional space, and the direction of view, which is set by three angles $(\alpha,\beta,\gamma)$. Thus, we set the trajectory of the camera motion using the dependencies $x=x(t)$, $y=y(t)$, $z=z(t)$, $\alpha=\alpha(t)$, $\beta=\beta(t)$, and $\gamma=\gamma(t)$. The camera’s first-person motion trajectory includes perspective projection operations with the camera initially fixed at the origin and the scene at a distance of $z_{\text{near}}$. Then, we apply transformations by rotating points around axes passing through the scene’s center and translating to this center. Acknowledging the limitations of a single-image-derived depth map, addressing distortions resulting from minor camera orientation deviations is crucial. Two essential considerations follow: adjusting scene position through infinitesimal transformations and employing the image-to-image approach after each transformation. The image-to-image technique facilitates the realization of seamless and semantically accurate transitions between frames, enhancing the aesthetic appeal of this approach. The examples of image-to-video generations are presented in the Image-to-Video appendix section.
@@ 6.4 Text-to-Video
Based on the Kandinsky 3.0 model, we also created the text-to-video generation pipeline Kandinsky Video [54], consisting, in addition to a text encoder and an image decoder, of two models – for keyframes generation and interpolation between them. Both models use the pretrained weights of Kandinsky 3.0 as the backbone. We have also implemented the publicly available text-to-video generation interface [^12] https://fusionbrain.ai/en/t2v/ in the Fusionbrain.ai website, which we mentioned above (Section 2). Please refer to the main paper for additional details regarding the text-to-video model.
@@ 7 Human evaluation
Some examples of text-to-image generations can be found in the Text-to-Image appendix section. To compare Kandinsky 3.0 and Kandinsky 3.1 with other well-known models, we have compiled a balanced set of 2.1K prompts in 21 categories. Using this set, we first performed several side-by-side (SBS) comparisons between different versions of the Kandinsky 3.0 and then selected the best version. We conducted three rounds of SBS comparisons involving 28 people to do this. Next, we conducted side-by-side comparisons of Kandinsky 3.0 with the Kandinsky 2.2 [11], SDXL [9] and DALL-E 3 [10] models. Each study involved 12 people who voted an average of 30,000 times in each SBS. For this purpose, we developed chatbot that showed one of 2.1K pairs of images.
@@
We compared the visual quality and text comprehension in total for all categories, and each category separately. The visual examples of Comparison to prior works and Human evaluation results for text-to-image generation are presented in the appendix.